For a given discrete non-negative random variable, 𝑝π‘₯ = 𝑃(𝑋 = π‘₯). Find π‘₯ such that 𝑝π‘₯ > 𝑝π‘₯+1, for the following distributions. a. 𝑋 is a binomial distribution with 𝑛 = 16 and 𝑝 = 0.5 b. 𝑋 is a Poisson distribution with πœ† = 8 c. 𝑋 is a Negative Binomial with π‘Ÿ = 5 and 𝑝 = 0.6 State the value of π‘₯ in the above sections. Verify whether the distributions are skewed for each section.


1
Expert's answer
2021-09-20T05:00:42-0400

(a) 𝑋 is a binomial distribution with 𝑛 = 16 and 𝑝 = 0.5.

"p_x=P(X=x)=\frac{n!}{x!(n-x)!}p^xq^{n-x}"

"\frac{p_{x+1}}{p_x}=\left(\frac{n!}{(x+1)!(n-x-1)!}p^{x+1}q^{n-x-1}\right) /\left( \frac{n!}{x!(n-x)!}p^xq^{n-x}\right)=\frac{n-x}{x+1}\frac{p}{q}\leq 1"

"p(n-x)\leq q(x+1)"

"pn-q\leq x(p+q)=x"

"pn-q=0.5\cdot16-0.5=7.5"

The minimal integer "x\geq 7.5" is 8.

Since "P(X=n-x)=\frac{n!}{x!(n-x)!}0.5^x0.5^{n-x}=\frac{n!2^{-n}}{x!(n-x)!}=P(X=x)", any binomial distribution with "p=0.5" is symmetric, hence, it is not skewed.

Answer. x=8


(b) 𝑋 is a Poisson distribution with πœ† = 8.

"p_x=P(X=x)=\frac{\lambda^{x}}{x!}e^{-\lambda}"

"\frac{p_{x+1}}{p_x}=\left(\frac{\lambda^{x+1}}{(x+1)!}e^{-\lambda}\right) /\left(\frac{\lambda^{x}}{x!}e^{-\lambda}\right)=\frac{\lambda}{x+1}\leq 1"

"x\geq \lambda-1=8-1=7"

The minimal integer "x\geq 7" is 7, this means that the mode of X is equal to 7.

"E(X)=\sum\limits_{x=0}^{\infty}xP(X=x)=\sum\limits_{x=0}^{\infty}x\frac{\lambda^{x}}{x!}e^{-\lambda}="

"=\lambda e^{-\lambda}\sum\limits_{x=1}^{\infty}\frac{\lambda^{x-1}}{(x-1)!}=\lambda e^{-\lambda}e^{\lambda}=\lambda=8"

Since "mode(X)=7<E(X)=8", the probability distribution is skewed to the left.

Answer. x=8, the probability distribution is skewed to the left.


(c) 𝑋 is a Negative Binomial with π‘Ÿ = 5 and 𝑝 = 0.6

"p_x=P(X=x)=\frac{(x+r-1)!}{x!(r-1)!}q^x p^r"

"\frac{p_{x+1}}{p_x}=\left(\frac{(x+r)!}{(x+1)!(r-1)!}q^{x+1} p^r\right) /\left(\frac{(x+r-1)!}{x!(r-1)!}q^x p^r\right)=\frac{x+r}{x+1}q\leq1"

"qx+qr\leq x+1"

"qr-1\leq px"

"x\geq (qr-1)/p=(0.4\cdot 5-1)/0.6=5/3"

The minimal integer "x\geq 5/3" is 2.

"E(X)=\sum\limits_{x=0}^{\infty}xP(X=x)=\sum\limits_{x=0}^{\infty}x\frac{(x+r-1)!}{x!(r-1)!}q^x p^r="

"=\frac{q}{p}r\sum\limits_{x=1}^{\infty}\frac{((x-1)+(r+1)-1)!}{(x-1)!((r+1)-1)!}q^{x-1} p^{r+1}=\frac{q}{p}r=\frac{0.4}{0.6}5=10/3"

Since "mode(X)=2<E(X)=10\/3", the probability distribution is skewed to the left.

Answer. x=2, the probability distribution is skewed to the left.